# MDCure: A Scalable Pipeline for Multi-Document Instruction-Following

### v1

@@ A.1 Source Texts
A.2 Data Generation Prompts
@@
To ensure a balanced composition for the pool of candidate instructions, we sample uniformly from all possible prompt templates during generation for each cluster of related documents. As mentioned in §3.1, once all instruction-answer pairs are generated, we append a brief direction indicating the expected length of the answer to the end of each instruction input (Figure 3). Options for this direction differ depending on whether an instruction was created using a General or Style-Specific template (Figure 9).
@@
A.3 Data Generation Hyperparameters
A.4 GPT Scoring Prompt
@@
A.5 MDCureRM Training Data Prompt
@@
A.6 MDCure Data Scoring
Given a candidate instruction-answer pair, MDCureRM generates a vector of six floating point values in the range $[0,1]$. To re-scale each value to the 1-5 range, we multiply each criterion score by 4 and add 1. To obtain the overall score for the instruction sample, we take a weighted sum of the resulting values, using a weight of $\frac{1}{9}$ for the general quality criteria (Relevance, Coherence & Factuality, Creativity) and a weight of $\frac{2}{9}$ for the multi-document specific criteria (Context Integration, Inter-Document Relationships, Complexity). In §5.5, we compare this weighting assignment with the evenly weighted alternative, which utilizes a weight of $\frac{1}{6}$ for each criterion in the case where all 6 criteria are used, and a weight of $\frac{1}{5}$ for each criterion in the case where one criterion is ablated at a time.
A.7 MDCure Dataset Statistics
@@
Appendix B Training Details
B.1 Instruction Tuning
@@
In all settings, we determine the best checkpoint for each model by comparing checkpoint performances in the zero-shot prompting setting on 500 test samples from MultiNews, Multi-XScience, HotpotQA-distractor, WikiHop, and QMDSCNN; we use LLM evaluation to assess MDS performance and F1 and EM scores to assess MHQA performance. Evaluation details are provided in App. C.4. For language modeling settings, we apply the loss to the instruction input as well during training, per recommendation by Shi et al. (2024).
B.2 MDCureRM Training
@@
Appendix C Evaluation Details
C.1 Benchmark Descriptions
@@ C.2 Baselines
C.3 Zero-Shot Evaluation Prompts
@@
C.4 LLM Evaluation Details
As noted in §4, to evaluate MDS performance, we turn to LLM evaluation to overcome limitations of $n$-gram-based methods such as ROUGE in long-context settings (Shaham et al., 2023). We employ the open-source LLM evaluation framework G-Eval (Liu et al., 2023), which is one of the most highly human-correlated evaluation frameworks to date when paired with GPT-4. For a given document and summary, G-Eval determines the quality of the summary by issuing four fine-grained scores for Relevance, Coherence, Consistency, and Fluency. The overall LLM score for a given summarization response is obtained by taking the average of the four scores. We adapt G-Eval to the multi-document setting using the prompts shown in Figures 14, 15, 16, and 17.
@@ C.4 LLM Evaluation Details
As GPT4 is expensive, we investigate use of several other commercial LLMs with the adapted G-Eval framework for evaluation on Multi-XScience and QMDSCNN, with the goal of ensuring strong alignment of resulting ratings with human evaluation. Based on computed correlations with scores from two human expert annotators on 500 samples per dataset, we find GPT-3.5-Turbo and Gemini-1.5-Flash to be most human-aligned for evaluation of model performances on Multi-XScience and QMDSCNN, respectively, and therefore use these for all reported evaluations.
Appendix D Meta-Evaluation on MDCureRM
@@
Appendix E Full Experimental Results
IT Data Setting HQA WH MXS QC SEAM ZeroScrolls (ZS) Avg
@@ Appendix E Full Experimental Results
Flan-T5 Base None 4.4 45.1 1.74 2.16 0.0 5.8 0.2 2.6 0.1 1.7 5.2 5.4 9.9 16.7 14.2 48.2 4.8 26.9 0.0 0.0 13.1 10.2
Unfiltered 31.7 47.5 4.04 2.35 0.0 6.7 0.1 2.7 0.2 1.9 5.3 8.6 12.4 15.5 28.7 46.6 6.2 30.0 54.3 2.4 21.0 16.1
GPT-Filtered 44.1 46.0 4.16 2.44 0.0 6.0 0.2 2.7 0.0 1.8 6.1 9.2 13.4 16.4 25.8 49.6 6.8 27.8 54.4 3.6 21.3 16.8
MDCureRM 47.3 48.3 4.22 2.58 0.0 7.5 0.3 2.7 0.0 2.1 6.6 9.2 14.5 16.1 34.6 48.2 8.1 31.2 54.4 3.4 22.6 17.9
Large None 24.4 54.6 3.19 2.78 1.1 7.9 0.1 3.1 0.3 2.5 19.5 13.5 15.5 9.4 24.4 31.0 21.1 11.8 44.4 41.1 23.2 17.3
Unfiltered 46.9 62.9 4.10 2.91 0.7 8.0 0.0 3.8 0.1 2.5 6.6 10.4 12.1 19.7 37.6 61.8 9.3 35.1 48.7 0.1 24.2 19.5
GPT-Filtered 46.3 65.1 4.13 2.88 0.5 8.3 0.0 3.9 0.1 2.6 7.2 9.9 11.9 19.8 39.1 63.4 8.3 33.8 48.8 0.0 24.2 19.7
MDCureRM 49.6 66.1 4.19 2.97 1.2 9.1 0.1 4.2 0.0 2.9 7.5 10.1 12.8 20.0 43.2 64.0 10.7 35.4 48.6 0.1 25.3 20.5
Qwen2-Ins 1.5B None 21.5 17.8 4.20 3.30 9.4 10.6 0.5 5.0 13.0 7.7 17.8 11.4 15.0 13.3 33.8 48.2 16.7 14.5 58.7 10.5 24.0 17.1
Unfiltered 32.9 30.5 4.24 3.57 15.7 12.0 0.4 5.0 16.8 10.0 14.1 9.8 15.3 15.7 40.3 48.0 15.7 16.1 53.8 10.1 23.9 18.9
GPT-Filtered 33.3 32.9 4.24 3.66 16.1 12.0 0.5 5.1 15.9 9.9 17.9 11.3 14.5 13.8 35.9 49.2 16.5 16.0 56.9 10.1 24.2 19.3
MDCureRM 37.7 34.8 4.25 3.73 16.4 12.0 0.6 5.9 18.1 10.6 19.8 12.0 14.9 13.1 38.0 55.0 16.4 19.7 54.9 12.5 25.6 20.5
7B None 30.5 39.6 4.30 3.57 5.0 11.9 0.5 6.4 13.1 7.4 21.1 12.2 15.0 13.2 33.8 72.8 19.0 29.4 16.3 6.5 23.9 18.6
Unfiltered 40.5 43.3 4.26 3.79 8.0 15.3 0.5 6.6 11.9 8.5 14.4 11.9 16.5 17.5 42.7 70.2 17.4 29.6 35.1 7.4 26.3 20.9
GPT-Filtered 42.0 44.0 4.26 3.84 8.7 15.3 0.9 6.6 12.1 8.7 11.5 12.0 16.8 18.1 43.1 67.4 15.9 28.4 54.1 19.6 28.7 22.3
MDCureRM 44.7 46.0 4.28 3.93 13.8 15.4 0.6 6.7 14.9 10.3 25.8 13.2 15.3 14.4 45.8 74.8 19.7 30.8 26.0 31.7 29.8 23.6
+PRIMERA 44.6 45.4 3.96 2.17 5.9 15.3 0.5 6.6 12.4 8.1 25.4 12.3 14.6 24.4 26.4 90.5 21.1 22.2 37.8 2.6 27.7 21.8
+QAMDen 42.6 45.8 3.92 2.37 9.5 15.3 0.5 6.4 12.4 8.8 18.8 4.5 15.7 36.0 41.5 76.2 13.6 18.4 50.7 9.5 28.5 22.3
Llama3.1-Ins 8B None 35.5 27.1 4.28 2.94 10.5 15.0 0.6 7.6 17.4 10.2 23.7 5.8 5.2 10.5 4.6 75.6 13.3 0.7 47.2 0.3 18.7 16.2
Unfiltered 37.6 34.4 3.81 4.07 15.3 16.3 0.5 7.8 18.9 11.8 20.2 14.1 17.4 21.9 52.3 50.0 19.4 25.3 32.6 32.1 28.6 22.3
GPT-Filtered 38.0 42.3 4.29 3.95 8.7 16.0 0.6 7.8 18.3 10.3 20.3 14.2 17.5 22.2 53.6 59.8 19.5 26.2 36.1 26.4 29.6 22.9
MDCureRM 44.7 43.7 4.29 4.22 16.3 16.4 0.6 7.9 18.5 11.9 19.9 14.5 17.6 22.6 52.2 68.2 19.2 27.5 34.3 33.4 30.9 24.5
70B None 53.9 38.1 4.28 3.97 25.1 21.9 0.6 6.1 11.5 13.0 21.9 14.9 18.1 25.5 45.5 66.7 23.4 42.5 61.4 43.8 36.4 27.9
Unfiltered 55.9 40.6 3.99 3.17 3.6 21.9 0.6 6.3 11.2 8.7 24.3 16.2 19.7 24.5 39.7 81.0 18.8 31.3 60.6 55.9 37.2 27.3
GPT-Filtered 57.4 41.2 3.97 3.37 4.9 22.0 0.7 6.4 10.7 8.9 24.5 15.6 17.8 21.3 41.0 81.0 18.5 30.3 69.2 54.4 37.4 27.6
MDCureRM 58.4 45.5 4.28 3.99 25.2 22.1 0.7 6.7 12.0 13.3 21.5 15.4 18.3 29.8 50.7 83.4 19.4 33.6 66.0 38.7 37.7 29.2
    Baselines PRIMERA 0.4 0.5 3.18 1.09 7.1 7.7 0.5 3.9 10.4 5.9 6.9 4.6 3.5 0.6 1.5 13.4 9.7 0.2 1.9 0.0 4.2 4.1
QAMDen 1.8 1.9 2.86 1.22 0.0 0.4 0.0 1.0 3.1 0.9 6.0 6.2 5.7 2.4 6.1 3.4 7.1 0.9 0.1 0.0 3.8 2.6
LongAlign-7B 10.4 14.3 4.15 3.75 11.5 16.5 0.0 4.1 16.8 9.8 19.5 13.5 15.5 9.4 24.4 31.0 21.1 11.8 44.4 41.1 23.2 16.5
ProLong-8B 43.6 34.3 3.86 2.46 9.5 17.4 0.4 10.7 18.0 11.2 25.6 15.2 17.6 31.6 32.2 66.2 23.2 8.6 51.8 50.5 32.2 24.3
PL+MDCure 45.8 45.1 3.93 3.40 15.3 17.3 0.5 10.5 18.9 12.5 24.6 13.9 15.9 30.7 28.2 66.7 21.6 21.7 51.4 47.1 32.2 25.4
Jamba 1.5 Mini 47.5 41.8 4.24 3.92 20.1 14.2 0.0 5.3 20.4 12.0 27.4 15.3 18.2 40.9 30.9 85.7 20.1 23.6 43.0 46.4 35.1 26.8
GPT-4o 57.5 50.0 4.50 4.25 8.9 18.6 0.3 14.2 28.9 14.1 25.7 14.8 15.9 35.3 58.8 88.6 20.7 49.2 83.5 5.6 39.8 30.8
Gemini 1.5 Pro 66.6 55.6 4.22 3.59 21.4 17.8 0.6 15.1 30.1 17.0 22.1 14.3 14.3 53.9 36.1 85.7 18.1 36.1 39.4 0.0 32.0 28.2
Table 10: Full evaluation of MDCure’d models versus baselines on 6 benchmarks in the zero-shot prompting setting. Dataset abbreviations are described in App. C.1.1.

### v3

@@ B.1 Source Texts
B.2 Data Generation Prompts
@@
To ensure a balanced composition for the pool of candidate instructions, we sample uniformly from all possible prompt templates during generation for each cluster of related documents.
B.3 Data Generation Hyperparameters
B.4 Instruction Format
We format the MD instruction inputs to place the generated instruction after the source documents, but alternate formats may be equally effective. Once all instruction-answer pairs are generated, since the expected output length of downstream MD tasks has a high variance (e.g., keyword based QA vs. long-form summarization), we append a brief direction indicating the expected length of the answer (i.e., expected number of words or sentences).101010 Our preliminary experiments demonstrated non-negligible performance gains on downstream MD tasks with the attachment of such length direction. Options for this direction differ depending on whether an instruction was created using a General or Style-Specific template (Figure 9). As shown in Fig. 3, the direction is appended to the end of the finalized instruction input, separated by a space.
@@ B.4 Instruction Format
B.5 Sample Instruction Data
We provide several qualitative examples demonstrating the diverse, complex, and open-ended nature of our MDCure data, formatted as prompt-response pairs (Table 11). Note that while each instruction input is formatted as discussed in App. B.4, in Table 11 we present only the generated prompt for each input and omit the context documents for easier visibility.
Instruction Prompt Answer
What is the standout feature of Samsung’s Galaxy S10 fingerprint sensor identified by the majority of the articles? Answer with a single word or brief phrase. Magical unlocking transition.
Describe the mechanism by which synthetic sandalwood stimulates hair growth. Give a response that is 3-4 words. Activates receptor OR2AT4.
Describe the implications of NASCAR’s new reduced horsepower package being used at the Auto Club 400 while considering the historical context of the race distance changes and past driver performances. Respond with 1-2 sentences. The implementation of the reduced horsepower package at the Auto Club 400 marks a significant shift in NASCAR’s strategy for intermediate tracks, aiming to enhance competitive balance; historically, the race has been shortened from 500 miles to 400 miles since 2011, impacting driver strategies and outcomes, as seen in past performances of drivers like Kevin Harvick and Kyle Busch.
Elaborate on the emotional and artistic significance of Adam Lambert’s song "Feel Something" as it reflects his personal journey and professional struggles, highlighting how this piece serves as a precursor to his upcoming album. Craft your response in 6 sentences. Adam Lambert’s song "Feel Something" signifies a pivotal moment in his life, as it encapsulates his journey of overcoming personal struggles and creative disillusionment. Written during a phase of introspection, the track addresses Lambert’s emotional experiences of feeling detached amidst the chaos of his career, illustrating his realization that the pursuit of love and connection could not be separated from the quest for personal happiness. The song represents both a cathartic release and an assertive step towards reclaiming his artistic vision, allowing Lambert to finally embrace the importance of self-care and vulnerability. This piece serves as the emotional cornerstone of his upcoming album, which aims to chronicle his commitment to taking ownership of his happiness and strength through music. The collaborative efforts with writers Benedict Cork and Josh Cumbee during its creation reflect a renewed artistic authenticity that Lambert seeks to reclaim. Thus, "Feel Something" not only marks a return to his musical roots but also signals a larger narrative of resilience, artistry, and self-discovery that will unfold in his forthcoming work.
Table 11: Representative instances of high-quality instruction-answer pairs, encompassing various MD task formulations and answer lengths, produced using our MDCure framework.
B.6 Diversity Analysis
To provide further insight regarding the syntactical and semantic diversity of our MDCure data, we follow the methodology used in Wang et al. (2023); Köksal et al. (2023) to analyze our size-12K MDCure dataset versus 12K samples from the training sets of HotpotQA-distractor and WikiHop, which are human-annotated MD benchmark datasets curated specifically for fine-tuning evaluation, along with 12K samples from the Tulu V2 SFT mixture (Rafailov et al., 2024), which is a mixture of human-annotated and synthetic data curated for general instruction tuning.
As in Wang et al. (2023), for each dataset, we parse the instruction in each training sample using Berkeley Neural Parser (Kitaev and Klein, 2018; Kitaev et al., 2019) and extract either the contained auxiliary and its dependent verb (e.g., “do hear”) or the contained verb and its direct object (e.g., “summarize events”). Additionally, as in Zhu et al. (2024), we analyze the relative complexity of instruction-response pairs using the Deita-complexity-scorer model (Liu et al., 2023a). We report statistics regarding unique verb-noun combinations and average complexity of each dataset in Table 12.
Dataset MDCure HotpotQA WikiHop Tulu V2
Unique Verbs 210 302 248 197
Unique Noun+Verb/ Aux-Verb Pairs 1269 1392 552 422
Avg Complexity 2.65 2.07 1.60 2.10
Top 5 Verbs describe, analyze, provide, summarize, reflect direct, have, take, play, write take, compose, use, get, make answer, solve, ask, modify, explain
Table 12: Data statistics according to instruction composition.
Versus human-annotated SFT and MD datasets, MDCure data covers a comparably wide range of tasks encompassing various instruction types, as indicated by its greater spread of unique verb+noun pairs, diverse and complex top 5 verbs, and highest-rated average complexity. These comparisons demonstrate the ability for MDCure prompts to yield generation of diverse MD instructions concerning tasks beyond those seen in typical SFT data, supporting the complementary nature of our instruction data and confirming the design of MDCure to reflect a broad array of task types and content.
B.7 GPT Scoring Prompt
@@
B.8 MDCureRM Training Data Prompt
@@
B.9 MDCure Data Scoring
Given a candidate instruction-answer pair, MDCureRM generates a vector of six floating point values in the range $[0,1]$. To re-scale each value to the 1-5 range, we multiply each criterion score by 4 and add 1. To obtain the overall score for the instruction sample, we take a weighted sum of the resulting values, using a weight of $\frac{1}{9}$ for the general quality criteria (Relevance, Coherence & Factuality, Creativity) and a weight of $\frac{2}{9}$ for the multi-document specific criteria (Context Integration, Inter-Document Relationships, Complexity). In §A.2, we compare this weighting assignment with the evenly weighted alternative, which utilizes a weight of $\frac{1}{6}$ for each criterion in the case where all 6 criteria are used, and a weight of $\frac{1}{5}$ for each criterion in the case where one criterion is ablated at a time.
B.10 MDCure Dataset Statistics
@@
B.11 Long-Context & Few-Shot Data Construction
For long-context and few-shot post-training experiments (§5.3), we construct instruction inputs by incorporating distractor articles or few-shot examples, respectively. Distractor articles are identified by embedding all NewSHead documents using the Sentence Transformers model all-distilroberta-v1, [^11] We choose this embedding model from https://www.sbert.net/docs/sentence_transformer/pretrained_models.html due to its strong performance, long context length, and compatibility with cosine similarity. and taking the most cosine-similar documents to the documents in the source cluster. We choose similar instead of random documents to increase task difficulty and encourage identification of relevant documents in a long context, as recent work finds using packing related texts during training can assist models to read and reason across text boundaries (Shi et al., 2023). For few-shot training, exemplars are chosen randomly, with each exemplar appearing once across the duration of training.
Appendix C Training Details
C.1 Instruction Tuning
@@
In all settings, we determine the best checkpoint for each model by comparing checkpoint performances in the zero-shot prompting setting on 500 test samples from MultiNews, Multi-XScience, HotpotQA-distractor, WikiHop, and QMDSCNN; we use LLM evaluation to assess MDS performance and F1 and EM scores to assess MHQA performance. Evaluation details are provided in App. D.4. For language modeling settings, we apply the loss to the instruction input as well during training, per recommendation by Shi et al. (2024).
C.2 MDCureRM Training
@@
C.3 PPO Training
For PPO training of LLAMA3.1-8B-Instruct (§5.5), we utilized 4 A100 80GB GPUs. Training occurred over 72K samples for a single epoch, with a global batch size of 64, learning rate of 1e-5, and LORA as described in §C.1. Additional hyperparameter search was performed over learning rates of 1e-4 and 1e-3. All other hyperparameter settings remained consistent with default values.
Appendix D Evaluation Details
D.1 Benchmark Descriptions
@@ D.2 Baselines
D.3 Zero-Shot Evaluation Prompts
@@
D.4 LLM Evaluation Details
As noted in §4, to evaluate MDS performance, we turn to LLM evaluation to overcome limitations of $n$-gram-based methods such as ROUGE in long-context settings (Shaham et al., 2023). We employ the open-source LLM evaluation framework G-Eval (Liu et al., 2023b), which is one of the most highly human-correlated evaluation frameworks to date when paired with GPT-4. For a given document and summary, G-Eval determines the quality of the summary by issuing four fine-grained scores for Relevance, Coherence, Consistency, and Fluency. The overall LLM score for a given summarization response is obtained by taking the average of the four scores and normalizing the result to the 0-100 range. We adapt G-Eval to the multi-document setting using the prompts shown in Figures 14, 15, 16, and 17.
@@ D.4 LLM Evaluation Details
As GPT4 is expensive, we investigate use of several other commercial LLMs with the adapted G-Eval framework for evaluation on Multi-XScience and QMDSCNN, with the goal of ensuring strong alignment of resulting ratings with human evaluation. To validate our choice of LLMs to evaluate MDS performance, we employ two expert annotators similarly to the procedure used in Fabbri et al. (2021). We observe strong alignment between annotators as indicated by the inter-annotator agreements presented in Table 15, measured via the Krippendorff’s alpha coefficient. Given the demanding nature of the annotation task and cost of human expert annotations, since the agreement is high additional annotations were deemed not necessary.
Multi-XScience QMDSCNN
Krippendorff $\alpha$ 0.91 0.94
Table 15: Inter-annotator agreements for human scores on summarization performance across six base models.
Based on computed correlations with scores from the annotators on 500 samples per dataset, we find GPT-3.5-Turbo and Gemini-1.5-Flash to be most human-aligned for evaluation of model performances on Multi-XScience and QMDSCNN, respectively, and therefore use these for all reported evaluations. The correlations are displayed in Table 16. These values are comparable with the human-LLM score correlations observed in leading LLM evaluation works such as G-Eval (Liu et al., 2023b), from which we adapted our LLM evaluation prompts.
Multi-XScience QMDSCNN
Pearson’s $r$ 0.54 0.58
Table 16: Correlations between LLM and average human scores of summarization performance across six base models. GPT-3.5-Turbo is used for evaluation of Multi-XScience performance and Gemini-1.5-Flash is used for evaluation of QMDSCNN performance.
Appendix E Meta-Evaluation on MDCureRM
@@
Appendix F Full Experimental Results
We display full experimental results, previously abbreviated in Table 1, in Table 18.
Model / IT Data Setting HQA WH MXS QC SEAM ZeroScrolls (ZS) Avg
@@ Appendix F Full Experimental Results
FlanT5 Base PRIMERA 0.4 0.5 70.7 24.2 7.1 7.7 0.5 3.9 10.4 5.9 6.9 4.6 3.5 0.6 1.5 13.4 9.7 0.2 1.9 0.0 4.2 8.8
QAMDen 1.8 1.9 63.6 27.1 0.0 0.4 0.0 1.0 3.1 0.9 6.0 6.2 5.7 2.4 6.1 3.4 7.1 0.9 0.1 0.0 3.8 7.2
None 4.4 45.1 38.7 48.0 0.0 5.8 0.2 2.6 0.1 1.7 5.2 5.4 9.9 16.7 14.2 48.2 4.8 26.9 0.0 0.0 13.1 14.5
Unfiltered 31.7 47.5 89.8 52.2 0.0 6.7 0.1 2.7 0.2 1.9 5.3 8.6 12.4 15.5 28.7 46.6 6.2 30.0 54.3 2.4 21.0 23.2
GPT-Filtered 44.1 46.0 92.4 54.2 0.0 6.0 0.2 2.7 0.0 1.8 6.1 9.2 13.4 16.4 25.8 49.6 6.8 27.8 54.4 3.6 21.3 24.1
MDCureRM 47.3 48.3 93.8 57.3 0.0 7.5 0.3 2.7 0.0 2.1 6.6 9.2 14.5 16.1 34.6 48.2 8.1 31.2 54.4 3.4 22.6 25.4
Large None 24.4 54.6 70.9 61.8 1.1 7.9 0.1 3.1 0.3 2.5 19.5 13.5 15.5 9.4 24.4 31.0 21.1 11.8 44.4 41.1 23.2 24.0
Unfiltered 46.9 62.9 91.1 64.7 0.7 8.0 0.0 3.8 0.1 2.5 6.6 10.4 12.1 19.7 37.6 61.8 9.3 35.1 48.7 0.1 24.2 27.3
GPT-Filtered 46.3 65.1 91.8 64.0 0.5 8.3 0.0 3.9 0.1 2.6 7.2 9.9 11.9 19.8 39.1 63.4 8.3 33.8 48.8 0.0 24.2 27.5
MDCureRM 49.6 66.1 93.1 66.0 1.2 9.1 0.1 4.2 0.0 2.9 7.5 10.1 12.8 20.0 43.2 64.0 10.7 35.4 48.6 0.1 25.3 28.5
Qwen2-Ins 1.5B None 21.5 17.8 93.3 73.3 9.4 10.6 0.5 5.0 13.0 7.7 17.8 11.4 15.0 13.3 33.8 48.2 16.7 14.5 58.7 10.5 24.0 25.5
Unfiltered 32.9 30.5 94.2 79.3 15.7 12.0 0.4 5.0 16.8 10.0 14.1 9.8 15.3 15.7 40.3 48.0 15.7 16.1 53.8 10.1 23.9 27.7
GPT-Filtered 33.3 32.9 94.2 81.3 16.1 12.0 0.5 5.1 15.9 9.9 17.9 11.3 14.5 13.8 35.9 49.2 16.5 16.0 56.9 10.1 24.2 28.1
MDCureRM 37.7 34.8 94.4 82.9 16.4 12.0 0.6 5.9 18.1 10.6 19.8 12.0 14.9 13.1 38.0 55.0 16.4 19.7 54.9 12.5 25.6 29.4
7B LongAlign-7B 10.4 14.3 92.2 83.3 11.5 16.5 0.0 4.1 16.8 9.8 19.5 13.5 15.5 9.4 24.4 31.0 21.1 11.8 44.4 41.1 23.2 25.3
None 30.5 39.6 95.6 79.3 5.0 11.9 0.5 6.4 13.1 7.4 21.1 12.2 15.0 13.2 33.8 72.8 19.0 29.4 16.3 6.5 23.9 27.4
Unfiltered 40.5 43.3 94.7 84.2 8.0 15.3 0.5 6.6 11.9 8.5 14.4 11.9 16.5 17.5 42.7 70.2 17.4 29.6 35.1 7.4 26.3 29.9
GPT-Filtered 42.0 44.0 94.7 85.3 8.7 15.3 0.9 6.6 12.1 8.7 11.5 12.0 16.8 18.1 43.1 67.4 15.9 28.4 54.1 19.6 28.7 31.4
MDCureRM 44.7 46.0 95.1 87.3 13.8 15.4 0.6 6.7 14.9 10.3 25.8 13.2 15.3 14.4 45.8 74.8 19.7 30.8 26.0 31.7 29.8 32.7
Llama3.1-Ins 8B ProLong-8B 43.6 34.3 85.8 54.7 9.5 17.4 0.4 10.7 18.0 11.2 23.8 16 17.5 23.5 39.9 70.6 22.6 19.2 54.6 48.5 33.6 32.1
None 35.5 27.1 95.1 65.3 10.5 15.0 0.6 7.6 17.4 10.2 23.7 5.8 5.2 10.5 4.6 75.6 13.3 0.7 47.2 0.3 18.7 24.3
Unfiltered 37.6 34.4 84.7 90.4 15.3 16.3 0.5 7.8 18.9 11.8 20.2 14.1 17.4 21.9 52.3 50.0 19.4 25.3 32.6 32.1 28.6 31.1
GPT-Filtered 38.0 42.3 95.3 87.8 8.7 16.0 0.6 7.8 18.3 10.3 20.3 14.2 17.5 22.2 53.6 59.8 19.5 26.2 36.1 26.4 29.6 32.1
MDCureRM 44.7 43.7 95.3 93.8 16.3 16.4 0.6 7.9 18.5 11.9 19.9 14.5 17.6 22.6 52.2 68.2 19.2 27.5 34.3 33.4 30.9 34.0
70B Jamba 1.5 Mini 47.5 41.8 94.2 87.1 20.1 14.2 0.0 5.3 20.4 12.0 21.1 15.1 17.3 22.2 47.3 84.4 20 28.1 59.1 26 34.1 35.3
None 53.9 38.1 95.1 88.2 25.1 21.9 0.6 6.1 11.5 13.0 21.9 14.9 18.1 25.5 45.5 66.7 23.4 42.5 61.4 43.8 36.4 37.1
Unfiltered 55.9 40.6 88.7 70.4 3.6 21.9 0.6 6.3 11.2 8.7 20.5 14.8 17 32.4 49.7 77.4 18 33.1 62.8 23.4 34.9 34.1
GPT-Filtered 57.4 41.2 88.2 74.9 4.9 22.0 0.7 6.4 10.7 8.9 21.5 15.3 18.3 30.1 50 83.8 19.3 32.7 65.2 38.7 37.5 35.9
MDCureRM 58.4 45.5 95.1 88.7 25.2 22.1 0.7 6.7 12.0 13.3 21.5 15.4 18.3 29.8 50.7 83.4 19.4 33.6 66.0 38.7 37.7 38.5
GPT-4o 57.5 50.0 100.0 94.4 8.9 18.6 0.3 14.2 28.9 14.1 25.7 14.8 15.9 35.3 58.8 88.6 20.7 49.2 83.5 5.6 39.8 40.6
Gemini 1.5 Pro 66.6 55.6 93.8 79.8 21.4 17.8 0.6 15.1 30.1 17.0 22.1 14.3 14.3 53.9 36.1 85.7 18.1 36.1 39.4 0.0 32.0 36.9
Table 18: Full evaluation of MDCure’d models versus baselines on 6 benchmarks in the zero-shot prompting setting. The rightmost “Avg” column reports the average of individual dataset scores. Dataset abbreviations are described in App. D.1.1. Rows specified by “MDCureRM” refer to our full MDCure pipeline applied to the corresponding base model and size. Size-comparable baselines are highlighted in blue and results of our final method is highlighted in yellow. Bold numbers show best performance in each group.